# Do Audio-Language Models Understand Linguistic Variations?

### v1

@@ Abstract
Open-vocabulary audio language models (ALMs), like Contrastive Language Audio Pretraining (CLAP), represent a promising new paradigm for audio-text retrieval using natural language queries. In this paper, for the first time, we perform controlled experiments on various benchmarks to show that existing ALMs struggle to generalize to linguistic variations in textual queries. To address this issue, we propose RobustCLAP, a novel and compute-efficient technique to learn audio-language representations agnostic to linguistic variations. Specifically, we reformulate the contrastive loss used in CLAP architectures by introducing a multi-view contrastive learning objective, where paraphrases are treated as different views of the same audio scene and use this for training. Our proposed approach improves the text-to-audio retrieval performance of CLAP by 0.8%-13% across benchmarks and enhances robustness to linguistic variation. Code and data will be open-sourced upon acceptance.
@@
Ramaneswaran Selvakumar1, Sonal Kumar1, Hemant Kumar Giri2, Nishit Anand1, Ashish Seth1, Sreyan Ghosh1, Dinesh Manocha1 1University of Maryland, College Park, 2NVIDIA, Bangalore {ramans, sonalkum, nishit, aseth125, sreyang, dmanocha}@umd.edu hgiri@nvidia.com
@@ 1 Introduction
    We present the first study to evaluate the robustness of ALMs for T2A retrieval. We construct five new benchmarks (synthetically with human-in-the-loop) to evaluate the performance of CLAP models in T2A retrieval across linguistically varied queries with similar intent. Our evaluation shows a consistent and significant drop in retrieval recall scores (0.1% - 16%) across our benchmarks, highlighting the vulnerability to linguistic variation.
@@
ML-ACT 35.53 34.87 50.68 48.14 21.52 17.91 08.72 06.06 10.12 08.77
MSCLAP-22 84.74 84.63 75.90 69.48 27.73 23.72 14.33 11.87 39.91 30.99
MSCLAP-23 80.77 77.63 79.71 71.10 55.12 39.15 38.27 24.89 47.84 39.21
CompA 97.17 96.23 79.22 73.85 43.03 40.24 33.32 23.56 49.54 39.51
@@ 5 Results And Analysis
Quantitative Analysis: Table 1 shows that current ALMs struggle with linguistic variations, as evidenced by a significant drop (0.1%-16%) in recall scores for paraphrased captions compared to the original captions. In contrast, RobustCLAP not only 1) improves recall scores on the original benchmarks by 0.8% to 13% compared to its base model but also 2) mitigates the performance drop on the paraphrased benchmarks, improving scores by 2% to 12% compared to the respective second best-performing model. CompA and MSCLAP-23, being trained on SoundDesc, perform better on that dataset. However, they show a significant 10-14% drop on the paraphrased SoundDesc benchmark, illustrating that fine-tuning can worsen the issue. We evaluate RobustCLAP on zero-shot audio classification tasks using ESC-50 Piczak (2015) and FSD50K Fonseca et al. (2022). CLAP gets a mAP@10 score of 94.25 and 52.20, while RobustCLAP gets 94.07 and 52.81, respectively, on ESC-50 and FSD-50K. We observe a negligible drop in performance, which indicates that prior knowledge is retained. RobustCLAP also outperforms ALMs on paraphrased audio-to-text retrieval, these results are indicated in Table 8. We also provide additional results on another CLAP variant in Table 7 to demonstrate the generalizability of our approach.
@@
Impact Of Sound Event And Attributes: In an acoustic scene, such as the "steady humming of an engine," the sound event refers to the sound and entity producing the sound (e.g., the "humming of an engine"); sound attributes describe its qualities (e.g., "steady"). We study how paraphrasing these elements affects retrieval performance by instructing the LLM to replace specifically the event and attributes with synonyms while maintaining the original linguistic structure. In Table 2 we observe that paraphrasing sound attributes leads to a 3.8% drop in Recall@10, while RobustCLAP significantly reduces this decline to just 0.4%. However, it is important to note that only 20% of the samples contain sound attributes, which limits the overall effect of this variation. Paraphrasing sound sources, on the other hand, has a much more significant impact, with recall dropping by as much as 15%. RobustCLAP mitigates this effect substantially, reducing the performance drop to 3%.
@@ F.2 Paraphrase Samples
AudioCaps CLAP 07.02 04.62 48.71 35.72 08.78 05.35 49.15 29.72
RobustCLAP 06.50 05.89 49.49 46.94 07.48 07.71 46.56 43.80
Clotho CLAP 38.08 30.90 81.24 75.88 37.03 30.72 81.91 74.83
RobustCLAP 38.27 37.12 81.53 80.76 39.43 37.32 82.49 82.30
Audioset SL CLAP 12.78 11.83 48.20 42.76 16.52 11.90 52.75 43.71
RobustCLAP 15.77 14.89 49.83 47.52 15.84 14.41 50.37 47.99
SoundDesc CLAP 04.95 03.66 21.91 16.95 03.23 02.34 17.75 13.63
RobustCLAP 05.45 04.12 23.05 20.83 03.78 02.95 19.08 16.92
DCASE CLAP 15.05 11.84 49.25 42.63 17.25 10.93 54.86 43.53
RobustCLAP 16.75 14.54 54.26 49.75 14.84 13.14 48.65 45.94
Table 7: T-A and A-T retrieval results for CLAP and RobustCLAP across the original test set (TEST) and the paraphrased test sets (TEST-P). We observe for this new variant of CLAP (630k-fusion-best) again across the benchmark there is a consistent and considerable drop in retrieval performance. When we derive RobustCLAP from this CLAP variant, we again notice the trend where there is improvement in general retrieval recall on TEST and mitigation of the drop in score in TEST-P. The best scores for each benchmark are highlighted in bold. This CLAP variant can be found here: https://huggingface.co/lukewys/laion_clap/blob/main/630k-fusion-best.pt
TEST TEST-P TEST TEST-P TEST TEST-P TEST TEST-P
@@
Clotho ML-ACT 12.87 11.42 50.68 48.14 13.20 12.03 52.71 48.87
MSCLAP-22 36.19 29.78 75.90 69.48 19.76 12.24 51.89 45.93
MSCLAP-23 37.03 30.47 79.71 71.10 22.87 16.26 61.53 51.19
CompA 36.39 29.11 79.22 73.85 17.14 11.97 53.44 44.34
LAION-CLAP 36.75 32.54 80.67 77.61 37.03 30.72 81.91 74.83
RobustCLAP 39.43 38.66 83.16 83.83 39.43 37.32 82.49 82.30

### v2

@@ Abstract
Open-vocabulary audio language models (ALMs), like Contrastive Language Audio Pretraining (CLAP), represent a promising new paradigm for audio-text retrieval using natural language queries. In this paper, for the first time, we perform controlled experiments on various benchmarks to show that existing ALMs struggle to generalize to linguistic variations in textual queries. To address this issue, we propose RobustCLAP, a novel and compute-efficient technique to learn audio-language representations agnostic to linguistic variations. Specifically, we reformulate the contrastive loss used in CLAP architectures by introducing a multi-view contrastive learning objective, where paraphrases are treated as different views of the same audio scene and use this for training. Our proposed approach improves the text-to-audio retrieval performance of CLAP by 0.8%-13% across benchmarks and enhances robustness to linguistic variation. We make our code publicly available  [^1] https://github.com/ramaneswaran/linguistic_robust_clapThese authors contributed equally to this work.
@@
Ramaneswaran Selvakumar*1, Sonal Kumar*1, Hemant Kumar Giri*2, Nishit Anand1, Ashish Seth1, Sreyan Ghosh1, Dinesh Manocha1 1University of Maryland, College Park, 2NVIDIA, Bangalore {ramans, sonalkum, nishit, aseth125, sreyang, dmanocha}@umd.edu hgiri@nvidia.com
@@ 1 Introduction
    We present the first study to evaluate the robustness of ALMs for T2A retrieval. We construct five new benchmarks (synthetically with human-in-the-loop) to evaluate the performance of CLAP models in T2A retrieval across linguistically varied queries with similar intent. Our evaluation shows a consistent drop in retrieval recall scores (0.1% - 16%) across our benchmarks, highlighting the vulnerability to linguistic variation.
@@
ML-ACT 35.53 34.87 27.54 23.90 21.52 17.91 08.72 06.06 10.12 08.77
MSCLAP-22 84.74 84.63 86.74 43.94 27.73 23.72 14.33 11.87 39.91 30.99
MSCLAP-23 80.77 77.63 51.14 42.12 55.12 39.15 38.27 24.89 47.84 39.21
CompA 97.17 96.23 51.28 42.49 43.03 40.24 33.32 23.56 49.54 39.51
@@ 2.1 Retrieval With Query Variations
2.2 Synthetic Data For Retrieval
Synthetic data generation has been widely studied in text-based representation learning and information retrieval. InPars Bonifacio et al. (2022), InParsv2 Jeronymo et al. (2023) and Promptagator Dai et al. (2022) generate synthetic queries from unlabelled documents for language encoder training. DINO Schick and Schütze (2021) generates synthetic textual similarity pairs for training cross-encoders while Gecko Lee et al. (2024) extensively uses LLMs to generate synthetic queries and hard-negatives. LARMOR Khramtsova et al. (2024) uses LLMs to generate synthetic data to adapt textual retrievers to a specific domain. On the other hand, synthetic data for improving audio-language models (ALMs) is still under explored. Approaches like CompA Ghosh et al. (2024b) pioneer the use of synthetic data to improve general and compositional representation of ALMs and train their models from scratch. In contrast, our approach adapts any off-the-shelf CLAP model and, with minimal additional training, enhances its robustness to linguistic variations while preserving its pre-trained knowledge and capabilities.
@@ 5 Results And Analysis
Quantitative Analysis: Table 1 shows that current ALMs struggle with linguistic variations, as evidenced by a significant drop (0.1%-16%) in recall scores for paraphrased captions compared to the original captions. In contrast, RobustCLAP not only 1) improves recall scores on the original benchmarks by 0.8% to 13% compared to its base model but also 2) mitigates the performance drop on the paraphrased benchmarks, improving scores by 2% to 12% compared to the respective second best-performing model. CompA and MSCLAP-23, being trained on SoundDesc, perform better on that dataset. However, they show a significant 10-14% drop on the paraphrased SoundDesc benchmark, illustrating that fine-tuning can worsen the issue. We evaluate RobustCLAP on zero-shot audio classification tasks using ESC-50 Piczak (2015) and FSD50K Fonseca et al. (2022). CLAP gets a mAP@10 score of 94.25 and 52.20, while RobustCLAP gets 94.07 and 52.81, respectively, on ESC-50 and FSD-50K. We observe a negligible drop in performance, which indicates that prior knowledge is retained. RobustCLAP also outperforms ALMs on paraphrased audio-to-text retrieval, these results are indicated in Table 7.
@@
Impact Of Sound Event And Attributes: In an acoustic scene, such as the "steady humming of an engine," the sound event refers to the sound and entity producing the sound (e.g., the "humming of an engine"); sound attributes describe its qualities (e.g., "steady"). We study how paraphrasing these elements affects retrieval performance by instructing the LLM to replace specifically the event and attributes with synonyms while maintaining the original linguistic structure. In Table 2 we observe that paraphrasing sound attributes leads to a 3.8% drop in Recall@1, while RobustCLAP significantly reduces this decline to just 0.4%. However, it is important to note that only 20% of the samples contain sound attributes, which limits the overall effect of this variation. Paraphrasing sound sources, on the other hand, has a much more significant impact, with recall dropping by as much as 15%. RobustCLAP mitigates this effect substantially, reducing the performance drop to 3%.
@@ 8 Limitations and Future Work
  • We use relatively shorter audio segments, in future this work can extended to long audio.
@@ F.2 Paraphrase Samples
Clotho ML-ACT 12.87 11.42 27.54 23.90 13.20 12.03 52.71 48.87
MSCLAP-22 36.19 29.78 86.74 43.94 19.76 12.24 51.89 45.93
MSCLAP-23 37.03 30.47 51.14 42.12 22.87 16.26 61.53 51.19
CompA 36.39 29.11 51.28 42.49 17.14 11.97 53.44 44.34
LAION-CLAP 36.75 32.54 52.03 43.98 37.03 30.72 81.91 74.83
RobustCLAP 39.43 38.66 57.27 53.48 39.43 37.32 82.49 82.30